Current Status of Each SOHO Instrument

Location: on McSoc: /Users/socfiles/info/current.html
To edit this file, edit the /u/ecs/soc/info/ copy

First Version:
15 February 2006, Zamkoff & Haugan

05 October 2007, Westenburger
08 March 2006, Zamkoff
09 March 2006, Zamkoff - final first version
09 March 2006, Zamkoff - CDS & UVCS QT increase due to UVCS current, LASCO QT increase for EIT bakeouts
24 March 2006, Zamkoff - make it html document and change document name
13 December 2006, Zamkoff - add SUMER detector A status and heaters on, MDI seeing SUMER temps, MDI on redundant TM/TC channel
17 July 2007, Zamkoff - change SWAN procedures for ESR
early October 2007, Gavin - added new TSTOLs info
16 October 2007, Zamkoff - change SWAN procedures with details on maneuvers, rolls
9 Jan 2009 EBZ - change GOLF heater 69 to be off
1 May 2009 EBZ - updated SUMER and LASCO
17 June 2009 EBZ - updated Exp LCL Mon note at bottom
22 June 2009 EBZ - Added note to UVCS about the LASCO debris issue
13 July 2009 EBZ - Added note about GOLF violation in Low Rate
30 July 2009 EBZ - Added note about C1 and FP for LASCO
25 Aug 2009 EBZ - Added note about ERNE BIAS2
09 Nov 2009 EBZ - Added note about new LASCO/EIT std mon channels

  • As of Jan 2009, only GOLF, UVCS, and LASCO will be turned off for Exp LCL Monitoring and filter set to 3 for a faster reaction time
  • ESR flag recipients: cds, celias, cepac, golf, lasco, mdi, sumer, swan, uvcs
  • Offpoint flag recipients: uvcs, lasco/eit, mdi, sumer
  • All instruments get OBT during Warm & Chilly Startups
  • When OBT sent only to certain instruments (such as in ESR and Thermal Reconfigurations), give to: MDI, SWAN, CEPAC, CELIAS, UVCS, LASCO (OCD#1834, 7/22/03)